# HyPlaneHead: Rethinking Tri-plane-like Representations in Full-Head Image Synthesis

### v1

@@ Appendix A Overview
This supplementary document provides additional materials to support the main paper "HyPlaneHead: Rethinking Tri-plane-like Representations in Full-Head Image Synthesis". It includes high-resolution qualitative comparisons, detailed explanations of the Near-Equal-Area Warping and Hy-Plane (2+2) representation. Additionally, we provide a video and key code snippets for reference.
@@
Appendix B High-Resolution Qualitative Comparison (Fig. 5)
Due to the page limit of the paper, we were only able to include a low-resolution version of the qualitative comparison (Fig. 5). To better demonstrate the advantages of our method in generating fine details, we provide a high-resolution version of Fig. 5.
fig.˜6 presents a high-resolution qualitative comparison with state-of-the-art methods. Each subfigure corresponds to the following representations: (a) Tri-plane representation from Chan et al. (2022). (b) Tri-grid representation from An et al. (2023). (c) Single spherical tri-plane representation from Li et al. (2024), where the white dashed box highlights a discontinuity in the hair region caused by seam artifacts. (d–e) Dual spherical tri-plane representation from Li et al. (2024). (f–j) Our proposed Hy-plane representation.
@@ Appendix C Details of Near-Equal-Area Warping
In the Lambert Azimuthal Equal-Area Projection, the south pole of the sphere is "opened" and then flattened into a circular domain centered at the north pole. During this unfolding process, the distances between latitude lines are adjusted such that the resulting circular projection maintains equal-area correspondence with the original spherical surface. A clearer understanding of this transformation can be gained from fig.˜7. fig.˜7(a–c) illustrates the dynamic process of the Lambert Azimuthal Equal-Area (LAEA) projection. fig.˜7(d) illustrates the Lambert Azimuthal Equal-Area (LAEA) projection using a world map example, demonstrating that it preserves area. Each orange circle represents a region of equal size on the original spherical surface.
Subsequently, we employ Elliptical Grid Mapping to convert the circular domain into a near-equal-area square grid. Among various methods for transforming a circle into a square, we choose Elliptical Grid Mapping due to its following advantageous properties: 1. Approximate equal-area mapping: The variation in local area across the transformed plane is minimized. 2. Smooth central region and minimal distortion at the boundaries: This preserves important structural details, especially near edges. 3.Computationally simple and stable: It avoids division operations, which is crucial for maintaining gradient stability during training. An intuitive illustration of this mapping is provided in fig.˜7. fig.˜7(e,f) show the deformation of Elliptical Grid Mapping under black-and-white stripe patterns, indicating that most regions experience minimal area distortion. fig.˜7(g,h) show the feature maps of the spherical plane before and after applying Elliptical Grid Mapping. Without this mapping, the model fails to effectively utilize the corner regions of the feature map. In contrast, with Elliptical Grid Mapping, most regions of the feature map are efficiently utilized.
@@
Figure 7: Illustrations of (a-d) the Lambert Azimuthal Equal-Area projection and (e-h) Elliptical Grid Mapping.
@@ Appendix D Details of Hy-Plane (2+2)
Appendix E More Qualitative Results
Below are additional qualitative results comparing our method with existing approaches on a larger set of examples. These include variations in pose, expression, and lighting conditions.

### v2

@@ Appendix A Overview
This supplementary document provides additional materials to support the main paper "HyPlaneHead: Rethinking Tri-plane-like Representations in Full-Head Image Synthesis". We present high-resolution qualitative comparisons that highlight the performance of our method against existing approaches in appendix˜B. A high-resolution qualitative comparison is provided in appendix˜C. We also provide a detailed explanation of the Near-Equal-Area Warping technique and our proposed Hy-Plane (2+2) representation, which are central to achieving view-consistent and artifact-free full-head image synthesis in appendix˜D. Comprehensive measurements of parameter count, training and inference speed, and VRAM usage across various methods and HyPlaneHead configurations are reported in appendix˜E. An in-depth discussion and comparison with related works, including OrthoPlanes, SYM3D, and other tri-plane-related algorithms, can be found in appendix˜F. Additional qualitative results across a broader set of examples are included to further demonstrate the effectiveness of our model in appendix˜G. We discuss the current limitations of our approach and potential directions for future work in appendix˜H. Finally, we include a section on Code of Ethics, where we address the ethical considerations and potential misuse of 3D head generation technologies in appendix˜I.
@@
Appendix B High-Resolution Qualitative Comparison (Main Paper Fig. 5)
Due to the page limit of the paper, we were only able to include a low-resolution version of the qualitative comparison (Main Paper Fig. 5). To better demonstrate the advantages of our method in generating fine details, we provide a high-resolution version of Main Paper Fig. 5.
fig.˜7 presents a high-resolution qualitative comparison with state-of-the-art methods. Each subfigure corresponds to the following representations: (a) Tri-plane representation from Chan et al. (2022). (b) Tri-grid representation from An et al. (2023). (c) Single spherical tri-plane representation from Li et al. (2024), where the white dashed box highlights a discontinuity in the hair region caused by seam artifacts. (d–e) Dual spherical tri-plane representation from Li et al. (2024). (f–j) Our proposed Hy-plane representation.
@@ Appendix C Details of Near-Equal-Area Warping
In the Lambert Azimuthal Equal-Area Projection, the south pole of the sphere is "opened" and then flattened into a circular domain centered at the north pole. During this unfolding process, the distances between latitude lines are adjusted such that the resulting circular projection maintains equal-area correspondence with the original spherical surface. A clearer understanding of this transformation can be gained from footnote˜2. footnote˜2(a–c) illustrates the dynamic process of the Lambert Azimuthal Equal-Area (LAEA) projection. footnote˜2(d) illustrates the Lambert Azimuthal Equal-Area (LAEA) projection using a world map example, demonstrating that it preserves area. Each orange circle represents a region of equal size on the original spherical surface.
Subsequently, we employ Elliptical Grid Mapping to convert the circular domain into a near-equal-area square grid. Among various methods for transforming a circle into a square, we choose Elliptical Grid Mapping due to its following advantageous properties: 1. Approximate equal-area mapping: The variation in local area across the transformed plane is minimized. 2. Smooth central region and minimal distortion at the boundaries: This preserves important structural details, especially near edges. 3. Computationally simple and stable: It avoids division operations, which is crucial for maintaining gradient stability during training. An intuitive illustration of this mapping is provided in footnote˜2. footnote˜2(e,f) show the deformation of Elliptical Grid Mapping under black-and-white stripe patterns, indicating that most regions experience minimal area distortion. footnote˜2(g,h) show the feature maps of the spherical plane before and after applying Elliptical Grid Mapping. Without this mapping, the model fails to effectively utilize the corner regions of the feature map. In contrast, with Elliptical Grid Mapping, most regions of the feature map are efficiently utilized.
@@
Figure 8: Illustrations of (a-d) the Lambert Azimuthal Equal-Area projection  [^1] https://en.wikipedia.org/wiki/Lambert_azimuthal_equal-area_projection and (e-h) Elliptical Grid Mapping  [^2] https://github.com/Kuuuube/Circular_Area/blob/main/wiki/mappings/elliptical_grid_mapping.md.
@@ Appendix D Details of Hy-Plane (2+2)
Appendix E Parameter Count, Speed and VRAM Usage Comparison
To provide a comprehensive comparison of parameter count, training and inference speed, as well as VRAM usage across different methods and HyPlaneHead configurations, we conducted measurements for each experiment listed in Table 1 of the paper. The results are shown in table˜2, where an asterisk (*) denotes that two spherical tri-planes are output simultaneously by a shared branch.
To provide a comprehensive and fair comparison, we clarify the definitions of the metrics used in our evaluation. Representation Parameters refer to the number of parameters in the feature maps of different tri-plane-like representations (e.g., tri-plane, spherical plane, hy-plane, etc.). Total Learnable Parameters denote the total number of trainable parameters in the entire model architecture, such as EG3D, PanoHead, SphereHead, and HyPlaneHead. Training Speed and Training VRAM Usage are measured on a single V100 GPU with a batch size of 2, representing the average time and memory consumption required to train 1,000 images. Similarly, Inference Speed and Inference VRAM Usage are evaluated under the same hardware setup but with a batch size of 1, by generating 100 images and computing the average time and memory cost per image.
From the statistics in the table, we can observe that different representations vary significantly in terms of feature plane parameter count. The tri-plane uses the fewest parameters (3 × 256 × 256 floating-point values), while the tri-grid with 512×512 resolution uses the most (9 × 512 × 512). Our proposed hy-plane uses 1 × 512 × 512 floating-point values, which is only 1.33 times the number used by the tri-plane. However, the overall difference in total learnable parameters across models is relatively small. The minor variations are mainly due to differences in the final convolutional layer configuration of StyleGAN2, which depends on how each representation is generated.
Notably, the total learnable parameters for entries No. 3 and No. 5 are identical. This is because in experiment No. 3, we did not modify the model code at all, which means only the rendering pipeline was adjusted to use one spherical tri-plane. We reported the actual parameter count to remain consistent with our experimental setup.
In terms of training and inference speed, as well as VRAM usage, there are some differences among the experiments. Overall, however, the additional computational overhead introduced by our innovations is relatively small.
Replacing the tri-plane with hy-plane introduces:
  • +5.5% training time,
  • +3.8% inference time,
  • +8.8% training VRAM,
  • -0.8% inference VRAM.
Adding the evenly split strategy further increases the cost by:
  • +8.4% training time,
  • +6.9% inference time,
  • +12.2% training VRAM,
  • +16.6% inference VRAM.
When combining with the near-equal-area projection, the overhead becomes:
  • +0.5% training time,
  • +3.5% inference time,
  • -5% training VRAM,
  • -5.7% inference VRAM.
Please note that in some cases, VRAM usage may actually decrease, likely due to memory fragmentation causing inaccuracies in the nvidia-smi measurement.
In summary, while our method does introduce some computational and memory overhead, the increase is relatively modest and justifiable given the significant improvements in disentanglement and reconstruction quality.
No. Representation Unify-Split Wrapping Rep. Params Total Params Training Speed (sec/kimg) Inference Speed (ms/image) Training VRAM (MiB) Infer. VRAM (MiB)
[…]
Table 2: Comparison of different 3D representation configurations in terms of parameter count, training/inference speed, and VRAM usage.
Appendix F Comparison and Discussion with Related Works
In this section, we provide a detailed comparison of our hy-plane representation with two closely related recent works: OrthoPlanes He et al. (2023) and SYM3D Yang et al. (2024), and discussion on compatibility with several tri-plane-related algorithms.
F.1 Comparison with OrthoPlanes
OrthoPlanes enhances the standard tri-plane representation by introducing multiple parallel planes along each Cartesian axis, effectively increasing the capacity of the feature field. This design is conceptually similar to the tri-grid formulation, where additional planar slices are used to capture finer geometric details.
However, this strategy does not address the underlying structural limitations of Cartesian-based projections. Specifically, OrthoPlanes still relies on axis-aligned planar projections in Cartesian Coordination, which inherently suffer from mirroring artifacts. As demonstrated in our main paper (fig.˜1), such artifacts manifest as mirroring-face artifacts and unnatural left-right duplication. Our hy-plane representation mitigates this by integrating spherical projection planes that naturally align with the radial symmetry of human heads, thereby disentangling symmetric and asymmetric components more effectively.
Moreover, the increased number of planes in OrthoPlanes leads to higher storage overhead. For downstream applications such as 3D head reconstruction or model initialization (e.g., in Portrait3D Wu et al. (2024b), AnimPortrait3D Wu et al. (2025), or ID-Sculpt Hao et al. (2025)), each sample must store K times more feature maps (where K is the number of parallel planes per axis), significantly increasing memory and bandwidth requirements. In contrast, our hy-plane uses only four planes (three planar + one spherical) while achieving superior fidelity.
Finally, due to the increased number of planes, it becomes difficult to integrate these approaches with our unify-split strategy. As a result, inter-channel feature penetration remains an issue in these methods.
F.2 Comparison with SYM3D
First, our goal differs from that of SYM3D. SYM3D enhances the symmetry of tri-plane representations through symmetric regularization, which is beneficial for generating fully symmetric artificial objects. In contrast, our hy-plane is designed to support a broader range of real-world scenarios where both symmetric and asymmetric structures coexist, such as in full-head portraits. Second, regarding feature penetration, SYM3D employs an attention-based scheme (View-wise Spatial Attention) to learn how to alleviate feature penetration across channels, whereas our hy-plane utilizes the unify-split strategy to geometrically and fundamentally prevent feature penetration at its source.
In fact, the correlation across $P_{XY}$, $P_{YZ}$ and $P_{XZ}$ planes discussed in SYM3D essentially corresponds to the inter-channel feature penetration problem we identify in our paper, though they observe it from a different angle. We detect this issue through visual inspection of feature maps (as shown in fig.˜1(a,b)), while SYM3D quantifies it using correlation metrics.
Inter-channel feature penetration causes similar values at the same UV positions across different feature maps, resulting in visually similar patterns, exactly what can be seen in our fig.˜1(a,b). This phenomenon is also visible in the top-right panel of fig. 7 in SYM3D, where GET3D’s feature maps exhibit repetitive vertical lines at the same spatial locations when zoomed in. Numerically, this leads to high correlations between different feature maps.
The difference lies in terminology: SYM3D does not explicitly refer to this as inter-channel feature penetration, but rather describes it as correlation. SYM3D attributes this issue to the use of single-view images during training, as opposed to multi-view data. While this is certainly true—more views provide richer geometric cues and reduce ambiguity—we also discuss this point in our paper. From a more fundamental perspective, however, the issue arises due to the structural nature of convolutional networks, where different channels are prone to interfere with each other, especially in the absence of direct supervision. Our solution addresses this root cause directly by modifying the architecture via the unify-split strategy, which effectively eliminates inter-channel feature penetration at its source.
In contrast, SYM3D uses view-wise spatial attention to alleviate the issue. As shown in the bottom-right panel of fig. 7 in SYM3D, this approach does reduce correlation to some extent. However, as illustrated in the middle-bottom panel, strong correlations still exist between certain planes, particularly $P_{YZ}$ and $P_{XZ}$. We have also computed and visualized similarity matrices similar to those in SYM3D’s fig. 7, and our results show significantly lower feature correlations across different planes.
Moreover, in terms of implementation complexity, our unify-split strategy is simpler and more intuitive compared to SYM3D’s view-wise attention mechanism, while achieving a more complete resolution of the problem.
F.3 Discussion on Compatibility with Tri-plane-related Algorithms
Dual Encoder Bilecen et al. (2024) introduces a dual-encoder GAN inversion approach for single-view 3D full-head reconstruction. It uses one encoder for the visible front region and another for the occluded back region, addressing issues such as mirroring artifacts in PanoHead’s W space. In contrast, our HyPlaneHead improves the underlying representation structure to reduce such artifacts and enhance W space quality. As a result, standard inversion methods like PTI—a general-purpose technique for common GANs—already yield better performance than previous approaches. We believe that combining our method with specialized inversion strategies like Dual Encoder, which are specifically tailored for full-head generation, could further improve results.
Tri2-plane Song et al. (2024) introduces a cascaded tri-plane representation across multiple scales of facial features, similar to a feature pyramid. This hierarchical design enables the model to generate both global structure and fine-grained details, resulting in richer and more detailed head reconstructions. Although it is still based on the tri-plane formulation and thus remains susceptible to mirroring artifacts, we believe that the multi-scale feature pyramid concept could be beneficially integrated with our hy-plane representation in future work.
Our method is also compatible with recent local editing approaches, such as Bilecen et al. (2025), which has been successfully applied to both tri-plane (EG3D) and tri-grid (PanoHead) representations. Given that our hy-plane shares a similar structure, we believe the method should be directly applicable to our representation as well.
Inspired by recent 3D avatar approaches Qian et al. (2024); Zhang et al. (2025); Qiu et al. (2025b); Kirschstein et al. (2025); Chu and Harada (2024); Qiu et al. (2025a); Kirschstein et al. (2024) based on 3D Gaussian Splatting (3DGS) Kerbl et al. (2023), we believe our hy-plane representation can be effectively integrated with 3DGS to further enhance rendering quality. We plan to explore this promising direction in future work.
Additionally, we believe that replacing the standard tri-plane representation in methods like LRM Hong et al. (2023) and InstantMesh Xu et al. (2024) with our hy-plane has the potential to improve their generation quality. On one hand, our unify-split strategy eliminates inter-plane feature penetration, allowing each plane to express its features more clearly and effectively. On the other hand, by incorporating a spherical plane, we enhance the model’s ability to represent asymmetric regions, such as facial details and hair on the back of the head. Therefore, we expect that integrating hy-plane into these models would lead to more accurate and artifact-free 3D reconstructions.
Appendix G More Qualitative Comparison
Figure 9 shows additional qualitative comparisons between our method and existing approaches on a larger set of examples. On the left side, samples (1–2) are generated by the official EG3D model, which is trained exclusively on the FFHQ dataset that lacks large-pose and back-view head images. As a result, these samples exhibit severe mirroring artifacts in their back views. Samples (3–4) use the tri-plane representation trained within our pipeline and data; while hair appears in the rear region, the front-facing facial attributes still dominate, indicating incomplete adaptation to non-frontal views. Samples (5–7) come from the official PanoHead model, which can produce detailed facial and hair textures but suffers from strong left-right symmetry and visible artifacts in the back view. Similarly, samples (8–10), using the tri-grid representation trained with our setup, also exhibit comparable symmetry and artifact issues. These problems—mirroring artifacts and severe left-right symmetry—are primarily caused by the Cartesian projection used in both tri-plane and tri-grid representations. Samples (11–13) are from the official SphereHead model, which addresses the mirroring issue through its spherical tri-plane design but results in more blurred outputs with less detailed facial and hair textures. The same trend is observed in samples (14–16), where a spherical tri-plane model is trained using our pipeline and data. In contrast, on the right side, samples (17–32) are generated by our HyPlaneHead model. By leveraging the novel hy-plane representation, our method not only eliminates mirroring artifacts but also achieves high-quality, consistent rendering from arbitrary view angles.
@@ Appendix G More Qualitative Comparison
Appendix H Limitations
Despite the promising results of our method, there are still several limitations that warrant further investigation. First, similar to previous works such as EG3D, PanoHead, and SphereHead, our model exhibits minor visual flickering or instability in fine details when rendering from gradually changing viewpoints. We believe this is partly due to the current GAN backbone’s limited capacity for high-fidelity view-consistent generation, and we plan to address this by adopting a more powerful generator architecture in future work.
Second, our method, like existing approaches, struggles with generating highly complex hairstyles such as ponytails, braids, or other structured hair arrangements. This limitation likely stems from insufficient training data covering such styles. We regard this as an important direction for future research and intend to expand our training dataset to include a broader variety of hairstyles and appearances.
Appendix I Code of Ethics
Our work presents a method for learning generalizable 3D full-head modeling from monocular images, which has potential applications in virtual avatars, digital content creation, and immersive experiences. However, such technology also raises ethical concerns, particularly regarding privacy and the potential for misuse, such as identity deception or unauthorized generation of realistic 3D head models. We are aware of these risks and emphasize the importance of responsible deployment, transparency, and user consent in any real-world application of this technology.